what are 2 consecutive even integers whose product plus two is 290
Answer by mananth(16949)

let the integers be x & (x+2)
x(x+2)+2=290
x^2+2x+2=290
Add -2 to both sides
x^2+2x-288=0
x^2+18x-16x-288=0
x(x+18)-16(x+18)=0
(x+18)(x-16)=0
x=-18, x=16
The numbers are 16 & 18
OR
-18, -16